Use the sample data and confidence level given below to complete parts​ (a) through​ (d). A research institute poll asked respondents if they felt vulnerable to identity theft. In the​ poll, n equals 1050 and x equals 569 who said​ "yes." Use a 95 % confidence level

​a) Find the best point estimate of the population proportion p.
Answer #1

​a) The best point estimate of the population proportion p

=\hat{p}=\frac{569}{1050}=0.5419

b)

Confidence interval formula is              
\hat{p}\pm Z*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}
=0.5419\pm 1.96*\sqrt{\frac{0.5419*(1-0.5419)}{1050}}
=(0.5118,0.5720)